# GS-ROR: 3D Gaussian Splatting for Reflective Object Relighting via SDF Priors

### v2

@@ 4.1. Deferred Gaussian splatting for relighting
4.2. SDF-aided Gaussian splatting
Thanks to the deferred Gaussian splatting, the blurry issue for reflective object relighting has been addressed. However, the geometry established by Gaussians exhibits erroneous surfaces and floaters, as it tends to overfit the appearance under training light conditions. Hence, more constraints on the geometry are needed. Inspired by NeRO and TensoSDF, we introduce SDF into our framework, as it has shown to be a good prior for geometry regularization. The key is leveraging the capability of SDF while avoiding the expensive rendering. For this, we first choose a performance-friendly backbone for SDF – TensoSDF. Then, we propose geometry-only supervision between Gaussians and SDF, or so-called mutual supervision, together with the color supervision between Gaussians renderings and the input images, without any SDF rendering. On the one hand, as Gaussians build the link between rendered images and input images, they can supervise the training of SDF. On the other hand, the SDF can also constrain the Gaussians.
@@ 4.2. SDF-aided Gaussian splatting
(9) $\mathcal{L}_{\rm sdf}=\lambda_{\rm gs2sdf}\mathcal{L}_{\rm gs2sdf}+\lambda_{\rm eik}\mathcal{L}_{\rm eik}+\lambda_{\rm hes}\mathcal{L}_{\rm hes}+\mathcal{L}_{\rm m}+\lambda_{\rm tv}\mathcal{L}_{\rm tv},$
@@ 4.3. SDF-aware pruning
where $p_{t}$ is an empirical hyper-parameter and set to $0.01$. In this way, the threshold is determined automatically and narrowed with the SDF in the optimization, so we avoid posing a constant too strict when the surface is under-reconstructed and prune the Gaussians too far away from the surface in the end.
@@ 5. Implementation details
The TensoSDF has a resolution of $400\times 400$ with feature channels set as $36$. The MLP decoder is a two-layer MLP whose width is 128. The Gaussian starts with 100K randomly initialized points.
Table 1. Losses used in our paper.
Name Apply on Description
$L_{\rm c}$ GS Color supervision for 3DGS
$L_{\rm nd}$ GS Normal Consistency loss from Gaussian Shader.
$L_{\rm sm}$ GS BRDF smoothness loss from Relightable Gaussian.
$L_{\rm\delta n}$ GS Normal regularization for Gaussian Shader.
$L_{\rm sdf2gs}$ GS The proposed depth/normal supervision for GS.
$L_{\rm gs2sdf}$ SDF The proposed depth/normal supervision for SDF.
$L_{\rm eik}$ SDF Eikonal loss from Gropp et al. (2020).
$L_{\rm hes}$ SDF Hession Loss from Zhang et al. (2022c).
$L_{\rm tv}$ SDF Total variance loss from TensoRF.
$L_{\rm m}$ GS/SDF Mask loss for both 3DGS and SDF.
@@
MII TensoIR TensoSDF NeRO GShader GS-IR RelightGS Ours
@@
We provide a table of loss functions (see  Tab. 1) for better understanding and more details of the losses are in the supplementary material.
We train our model on an NVIDIA RTX 4090 and optimize it with Adam optimizer (Kingma and Ba, 2017). The learning rate for the TensoSDF (Li et al., 2024) grid and the MLP decoder is $1e^{-2}$ and $1e^{-3}$, respectively. The learning rate of Gaussian components follows the setting in 3DGS (Kerbl et al., 2023), and all attributes not in vanilla 3DGS have a learning rate of $2e^{-3}$. The initial resolution of TensoSDF is set to $128$ and increases to the final resolution $400\times 400$, which uses only 60% capacity of the original TensoSDF, whose final resolution is $512\times 512$. After training, the TensoSDF will be discarded, and only Gaussians will be used for relighting. The pruning threshold is set to $0.01$ empirically. The optimization procedure takes about 1.5 hours in total.
@@ 6.1. Evaluation setup
We select 7 representative methods for comparison. We choose MII (Zhang et al., 2022b), TensoIR (Jin et al., 2023), TensoSDF (Li et al., 2024), and NeRO (Liu et al., 2023) for NeRF-based methods and Gaussian Shader (Jiang et al., 2024), GS-IR (Liang et al., 2024), and Relightable Gaussian (Gao et al., 2024) for Gaussian-based methods. We trained these models based on their public codes and configurations. Note that we add the mask loss to train all the compared methods, for stabilizing training and fair comparison.
@@
We use the peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) (Wang et al., 2004) to measure the quality of results for the comparison and ablation study in the main paper. We present the learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018) in the supplementary materials. Due to the ambiguity between albedo and light, we normalize the relighting images to match the average colors of ground-truth images before computing the metrics, following NeRO. When the ground truth albedo is available, we rescale the relighting images based on the ground truth albedo and predicted albedo, following TensoIR. We record the mean training time and the frame per second (FPS) [^1] NeRO and TensoSDF use the Cycles Render Engine in Blender to provide relighting results, so the FPS depends on samples per pixel, which is 1024 following the NeRO setting. in the evaluation, which tests on a single RTX 4090.
@@ 6.2. Comparison with previous works
We evaluate the relighting performance of specular materials on the Glossy Blender dataset. The quantitative measurements are shown in Tab. 2, and the relighting visualizations are placed in Fig. 10. Our method outperforms other Gaussian-based methods and achieves competitive results of the NeRF-based methods. We also show the decomposed maps in Fig. 11 and Fig. 12. Due to the SDF priors and the deferred Gaussian pipeline, our method can produce a smooth surface without losing details and predict reasonable BRDF parameters.
@@
Table 3. Relighting quality in terms of PSNR↑ and SSIM↑ on the TensoIR synthetic dataset. Numbers in bold indicate the best performance, and underscored numbers indicate the second best.
GShader GS-IR RelightGS Ours
@@
Table 4. NVS quality in terms of PSNR↑ and SSIM↑ on the Glossy Blender dataset. Numbers in bold indicate the best performance, and underscored numbers indicate the second best.
GS-IR RelightGS GShader Ours
@@
Figure 7. Ablation of several key components in our method, including the deferred shading, SDF supervision and pruning. The insets are shown in blue, and some floaters are shown in green.
@@
We also present the NVS results on the Glossy Blender dataset and the Shiny Blender Dataset. The quantitative evaluation in terms of PSNR and SSIM is present in Tabs. 4 and 5. Our method can provide high-quality NVS results with sharper highlights, which are present in Figs. 14 and 15, revealing the wide applications of our method. Our method outperforms the existing Gaussian-based methods on most scenes. However, our method fails in the Coffee scene, as the deferred pipeline assumes an opaque surface, and the liquid in the Coffee does not meet the assumption. Besides, we present a comparison of normal estimation in Fig. 16 and the mean angular error (MAE) in Tab. 6. Thanks to the SDF priors, our method can provide robust normal estimation, while the results of other Gaussian-based methods are overfitted or noisy.
@@ 6.3. Ablation study
Deferred Gaussian pipeline
After applying the deferred Gaussian pipeline, this variant provides a more specular result. However, due to the erroneous normal, the surface reflects light from the wrong directions, shown in the 2nd column of Fig. 7. We obverse more artifacts from the deferred results, for the pixel color from deferred shading corresponds to one shading point on the estimated surface while the one from forward shading is a combination of rendered colors of multiple Gaussians. The deferred formulation leaves less freedom for individual Gaussians to overfit. Consequently, more Gaussian outliers are generated to overfit sharp highlights, causing additional artifacts. Nevertheless, the deferred shading is necessary for specular materials (Sec. 4.1), and its issue has been solved in our full model.
@@
Then, we incorporate the TensoSDF and utilize its priors to regularize the geometry. The model without TensoSDF is vulnerable to the local minima and overfits under the training light condition, leading to erroneous geometry estimation. After employing the TensoSDF (3rd column in Fig. 7), the model mitigates the ambiguity between material and geometry and predicts a reasonable normal. We also validate the effectiveness of our mutual supervision by comparing it to supervision by a pretrained TensoSDF. As shown in Tab. 7, although using the pretrained TensoSDF for supervision could improve the quality slightly, it needs 5 hours for TensoSDF training. In contrast, our method achieves a balance between performance and efficiency. Besides, our method preserves the details while SDF-based methods fail, as shown in Fig. 8 (see the cat whisker), revealing the benefits of our Gaussian representation.
@@
Table 7. Comparison of the relighting quality between mutual supervision (Ours) and using a pretrained TensoSDF to supervise Gaussians unidirectionally on the Glossy Blender dataset. Despite the slight quality improvement, using the pretrained model leads to an extensive training time cost.
@@
Table 8. Ablation study of three key components on the Glossy Blender dataset. “Def.” means the deferred Gaussian pipeline, “Inc.” means the incorporation of SDF and Gaussian, and “Pru.” means the pruning operation.
@@ 6.4. Discussion and limitations
Our model improves the relighting quality compared to other Gaussian-based methods. However, some issues remain to be solved. The SDF representation used in our framework is used as a prior, and it can not be used for exporting a faithful mesh due to its low resolution. Besides, we obverse the split-sum approximation in DiffRast (Laine et al., 2020) causes blurry renderings for extremely low-roughness surfaces. We render a mirror-like plane using the Cycle path-tracing renderer in Blender and the split-sum approximation in the DiffRast. The comparison is in Fig. 9, and the split-sum result is blurry. We believe replacing it with a more accurate rendering method will improve the relighting quality. Last, we only consider direct lighting in our framework, and introducing indirect illumination will benefit the relighting quality. We leave it for future work.
@@
Figure 10. Relighting results on the Glossy Blender Dataset. Relightable Gaussian, GS-IR, TensoIR, and MII fail to model specular highlights. Gaussian Shader provides specular relighting results with obvious artifacts. Our method, along with NeRO and TensoSDF, provides high-quality results, while our method only uses 25% training time of TensoSDF and enables real-time relighting. We present the metrics per scene (PSNR/SSIM).
@@
Figure 13. Relighting results on real data from NeILF++. Relightable Gaussian and GS-IR fail to reconstruct the reflective objects in real scenes. Gaussian Shader cannot relight the specular highlights faithfully, while our method can provide reasonable relighting results.

### v3

@@ 4.1. Deferred Gaussian splatting for relighting
4.2. An rendering-free mutual supervision
Thanks to the deferred Gaussian splatting, the blurry issue for reflective object relighting has been addressed. However, the geometry established by Gaussians exhibits erroneous surfaces and floaters, as it tends to overfit the appearance under training light conditions. Hence, more constraints on the geometry are needed. Inspired by NeRO and TensoSDF, we introduce SDF into our framework, as it has been shown to be a good prior for geometry regularization. The key is leveraging the capability of SDF while avoiding the expensive rendering. For this, we first choose a performance-friendly backbone for SDF – TensoSDF. Then, we propose geometry-only supervision between Gaussians and SDF, or so-called mutual supervision, together with the color supervision between Gaussians renderings and the input images, without any SDF rendering. On the one hand, as Gaussians build the link between rendered images and input images, they can supervise the training of SDF. On the other hand, the SDF can also constrain the Gaussians.
@@ 4.2. An rendering-free mutual supervision
(9) $\mathcal{L}_{\rm sdf}=\lambda_{\rm gs2sdf}\mathcal{L}_{\rm gs2sdf}+\lambda_{\rm eik}\mathcal{L}_{\rm eik}+\lambda_{\rm hes}\mathcal{L}_{\rm hes}+\lambda_{\rm m}\mathcal{L}_{\rm m}+\lambda_{\rm tv}\mathcal{L}_{\rm tv},$
@@ 4.3. SDF-aware pruning
where $p_{t}$ is an empirical hyperparameter and set to $0.01$. In this way, the threshold is determined automatically and narrowed with the SDF in the optimization, so we avoid posing a constant too strict when the surface is under-reconstructed and prune the Gaussians too far away from the surface in the end.
4.4. GS-guided SDF enhancement
Figure 7. Although the relighting framework ensures smooth blended normal to render realistic glossy highlight, the blended depth is inconsistent with the normal, and thus the mesh of TSDF fusion from depth can be problematic.
Figure 8. The overview of GS-guided SDF enhancement. The TensoSDF learned from the first step is upsampled and refined by the blended normal from fixed Gaussians.
Due to the inconsistency between the normal and depth from blended Gaussians, as shown in Fig. 7, the meshes of TSDF fusion from depth can be problematic, even though the normal is well optimized in the GS framework. Lots of works (Yu et al., 2024b; Radl et al., 2024; Chen et al., 2024; Moenne-Loccoz et al., 2024) try to alleviate this inconsistency, while these solutions complicate the rendering pipeline and thus reduce the GS efficiency in training and rendering.
Since our framework has the TensoSDF as an extra geometry representation, we can directly extract mesh from it, without complicating the GS pipeline. Therefore, we propose the GS-guided SDF enhancement, which utilizes the normal from well-optimized Gaussians to enable extracting high-quality meshes from TensoSDF.
Specifically, we first upsample the TensoSDF to capture the fine-grained details, which keeps a low resolution for efficiency in the first step. Then, the upsampled TensoSDF is finetuned efficiently via the learned normal from fixed Gaussians, as shown in Fig. 8. The loss for finetuning $\mathcal{L}_{\rm ft}$ is
(12) $\mathcal{L}_{\rm ft}=(1-<\textbf{sg}[\textbf{n}_{\rm gs}],\textbf{n}_{\rm sdf}>)+\lambda_{\rm hes}\mathcal{L}_{\rm hes}+\lambda_{m}\mathcal{L}_{\rm m}+\lambda_{\rm tv}\mathcal{L}_{\rm tv},$
where all losses $\mathcal{L}_{[\cdot]}$ share the same definition as in Eqn. (9).
This way, the TensoSDF learns geometry details from the well-optimized Gaussians. Therefore, our method achieves high-quality geometry reconstruction for reflective objects, which beats the TensoSDF optimized via ray marching and uses only 33% training time of it. Compared to optimizing a full-resolution SDF and Gaussians jointly as in GSDF, we modulate the resolution of TensoSDF adaptively, leading to a more efficient framework and reducing the peak memory usage during optimization. Consequently, our framework maintains the highly efficient training/rendering from Gaussian splatting and meanwhile benefits from the robustness of the TensoSDF representation for geometry reconstruction.
@@ 5. Implementation details
The TensoSDF has a resolution of $400\times 400$ with feature channels set as $36$, which upsamples to $512\times 512$ further. The MLP decoder is a two-layer MLP whose width is 128. The Gaussian starts with 100K randomly initialized points.
@@
MII TensoIR TensoSDF NeRO GShader GS-IR R3DG Ours
@@
Table 2. Losses used in our paper.
Name Apply on Description
$L_{\rm c}$ GS Color supervision for 3DGS
$L_{\rm nd}$ GS Normal Consistency loss from GShader.
$L_{\rm sm}$ GS BRDF smoothness loss from R3DG.
$L_{\rm\delta n}$ GS Normal regularization for GShader.
$L_{\rm sdf2gs}$ GS The proposed depth/normal supervision for GS.
$L_{\rm gs2sdf}$ SDF The proposed depth/normal supervision for SDF.
$L_{\rm eik}$ SDF Eikonal loss from Gropp et al. (2020).
$L_{\rm hes}$ SDF Hession Loss from Zhang et al. (2022c).
$L_{\rm tv}$ SDF Total variance loss from TensoRF.
$L_{\rm m}$ GS/SDF Mask loss for both 3DGS and SDF.
@@
We provide a table of loss functions (see  Tab. 2) for better understanding and more details are in the supplementary material.
We train our model on an NVIDIA RTX 4090 and optimize it with Adam optimizer (Kingma and Ba, 2017). The learning rate for the TensoSDF (Li et al., 2024) grid and the MLP decoder is $1e^{-2}$ and $1e^{-3}$, respectively. The learning rate of Gaussian components follows the setting in 3DGS (Kerbl et al., 2023), and all attributes not in vanilla 3DGS have a learning rate of $2e^{-3}$. The initial resolution of TensoSDF is set to $128$ and increases to the final resolution $400\times 400$, which uses only 60% capacity of the original TensoSDF, whose final resolution is $512\times 512$. The optimization procedure takes about 1.5 hours in total. After training, only Gaussians will be used for relighting. To extract high-quality meshes from TensoSDF, a 15-minute finetuning is needed, using the Adam optimizer and a cosine learning rate schedule. The learning rate for the TensoSDF (Li et al., 2024) grid and the MLP decoder starts with $1e^{-2}$ and $1e^{-3}$, which decays to $5e^{-4}$ and $5e^{-5}$ finally.
@@ 6.1. Evaluation setup
We select 7 representative methods for comparison. We choose MII (Zhang et al., 2022b), TensoIR (Jin et al., 2023), TensoSDF (Li et al., 2024), and NeRO (Liu et al., 2023) for NeRF-based methods and GShader (Jiang et al., 2024), GS-IR (Liang et al., 2024), and R3DG (Gao et al., 2024) for Gaussian-based methods. While evaluating the surface quality, we also select some non-relightable Gaussian-based reconstruction methods, including 2DGS, GOF (Yu et al., 2024b), and PGSR (Chen et al., 2024). We trained these models based on their public codes and configurations. Note that we add the mask loss to train all the compared methods, for stabilizing training and fair comparison.
[…]
Non-relightable Relightable
2DGS GOF PGSR GShader GS-IR R3DG Ours
Angel 0.93 0.73 0.54 0.85 1.77 0.98 0.41
Bell 3.35 3.07 1.65 1.10 11.53 4.18 0.31
Cat 2.74 2.25 2.93 2.56 5.88 3.39 1.34
Horse 1.18 1.07 0.73 0.73 1.96 1.35 0.34
Luyu 1.35 1.46 0.85 1.07 2.25 1.68 0.81
Potion 4.50 4.13 1.96 4.74 6.23 3.80 0.75
Tbell 5.50 4.61 4.72 5.74 10.21 5.00 0.55
Teapot 2.18 3.32 1.24 3.40 7.19 4.79 0.47
Mean 2.72 2.58 2.44 2.53 5.88 3.15 0.62
@@
We use the peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) (Wang et al., 2004) to measure the relighting quality of results for the comparison and ablation study in the main paper. We present the learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018) in the supplementary materials. Due to the ambiguity between albedo and light, we normalize the relighting images to match the average colors of ground-truth images before computing the metrics, following NeRO. When the ground truth albedo is available, we rescale the relighting images based on the ground truth albedo and predicted albedo, following TensoIR. We use the mean angular error (MAE) and the Chamber distance (CD) to measure the quality of geometry reconstruction. We record the mean training time and the frame per second (FPS) [^1] NeRO and TensoSDF use the Cycles Render Engine in Blender to provide relighting results, so the FPS depends on samples per pixel, which is 1024 following the NeRO setting. in the evaluation. All the FPS and training time are tested on an RTX 4090 unless otherwise specified.
@@ 6.2. Comparison with previous works
We evaluate the relighting performance of specular materials on the Glossy Blender dataset. The quantitative measurements are shown in Tab. 1, and the relighting visualizations are placed in Fig. 14. Our method outperforms other Gaussian-based methods and achieves competitive results of the NeRF-based methods. We also show the decomposed maps in Fig. 15, Fig. 16, and Fig. 17. Due to the SDF priors and the deferred Gaussian splatting pipeline, our method can produce a smooth surface without losing details and predict reasonable BRDF parameters.
@@
[…]
GShader GS-IR R3DG Ours
@@
Table 5. NVS quality in terms of PSNR↑ and SSIM↑ on the Glossy Blender dataset. Numbers in red indicate the best performance, and numbers in orange indicate the second best.
GS-IR R3DG GShader Ours
@@
Figure 9. Ablation of several key components in our method, including deferred shading, SDF supervision, and pruning. The first, second, and third rows show relighting results, normal, and roughness, respectively. The insets are shown in blue, and some floaters are shown in green.
@@
We also present the NVS results on the Glossy Blender dataset and the Shiny Blender Dataset. The quantitative evaluation in terms of PSNR and SSIM is present in Tabs. 5 and 6. Our method can provide high-quality NVS results with sharper highlights, which are present in Figs. 19 and 20, revealing the wide applications of our method. Our method outperforms the existing Gaussian-based methods on most scenes. However, our method fails in the Coffee scene, as the deferred pipeline assumes an opaque surface, and the liquid in the Coffee does not meet the assumption.
Geometry quality
We present the comparisons of geometry quality on the Shiny Blender and Glossy Blender datasets. The choice of measurement is based on the available ground truth. The normal estimation on the Shiny Blender dataset in Fig. 21 and MAE in Tab. 7. Thanks to the SDF priors, our method can provide robust normal estimation, while the results of other Gaussian-based methods are overfitted or noisy.
The CD on the Glossy Blender dataset is in Tab. 3. Our method achieves the best reconstruction quality regarding CD among the Gaussian-based methods. These Gaussian-based methods have no specific design for reflective objects, and some methods (e.g., PGSR) assume multi-view consistency of appearance, which does not hold for reflective objects. Hence, all these methods perform poorly on the Glossy dataset. Although GSDF is most relevant to our method, it does not converge for reflective objects even with mask loss, and thus its result is not available. Besides, we show the mesh visualizations of representative methods in Fig. 22. Visually, the existing Gaussian-based methods cannot provide smooth surfaces. TensoSDF and NeRO can provide smooth meshes while discarding some geometry details (see the mane of ‘horse’). Our method can provide globally smooth meshes with geometry details.
@@ 6.3. Ablation study
Deferred Gaussian splatting pipeline
After applying the deferred Gaussian splatting pipeline, this variant provides a more specular result. However, due to the erroneous normal, the surface reflects light from the wrong directions, shown in the 2nd column of Fig. 9. We obverse more artifacts from the deferred results, for the pixel color from deferred shading corresponds to one shading point on the estimated surface while the one from forward shading is a combination of rendered colors of multiple Gaussians. The deferred formulation leaves less freedom for individual Gaussians to overfit. Consequently, more Gaussian outliers are generated to overfit sharp highlights, causing additional artifacts. Nevertheless, the deferred shading is necessary for specular materials (Sec. 4.1), and its issue has been solved in our full model.
@@
Then, we incorporate the TensoSDF and utilize its priors to regularize the geometry. The model without TensoSDF is vulnerable to local minima and overfits under the training light condition, leading to erroneous geometry estimation. After employing the TensoSDF (3rd column in Fig. 9), the model mitigates the ambiguity between material and geometry, thus predicting reasonable normal and material parameters (e.g., roughness). We validate the effectiveness of our mutual supervision by comparing it to supervision by a pretrained TensoSDF. As shown in Tab. 8, although using the pretrained TensoSDF for supervision could improve the quality slightly, it needs 5 hours for TensoSDF training. Optimizing GS with full-resolution SDF, as in GSDF, also improves the performance, increasing 40% (+8.5GB) memory usage. In contrast, our method achieves a balance between performance and efficiency. Besides, our method preserves the details while SDF-based methods fail, as shown in Fig. 10 (see the cat whisker), revealing the benefits of our Gaussian representation.
@@
Table 8. Comparison of the relighting quality between mutual supervision (Ours), using a pretrained TensoSDF to supervise Gaussians unidirectionally (Pretrained), and jointly optimization of full-resolution SDF with GS (Joint opt.) as in GSDF on the Glossy Blender dataset. Despite the slight improvement, using the pretrained model leads to an extensive training time or memory cost.
@@
Joint opt. 21.68 (+0.87/+8.5GB) 26.50 (+0.22/+8.5GB) 24.34 (+1.03/+8.5GB)
Figure 11. The reconstruction results from TSDF fusion, low-res. TensoSDF and finetuned TensoSDF. The meshes from TSDF fusion and low-res. TensoSDF has obvious artifacts, while our finetuned TensoSDF provides globally smooth meshes with fine-grained details.
@@
GS-guided TensoSDF enhancement
The above designs ensure a promising normal for relighting, while the enhancement enables high-quality geometry reconstruction. We show the visualization of meshes from TSDF fusion, low-resolution TensoSDF, and finetuned TensoSDF in Fig. 11. Due to the inconsistency between the normal and depth of Gaussian primitives, though the blended normal is smooth, the artifacts are apparent in the mesh from TSDF fusion. Furthermore, the low-resolution TensoSDF cannot provide smooth meshes and leads to grid-like artifacts. After finetuning, the upsampled TensoSDF provides globally smooth meshes with fine-grained details. We also measure the mesh quality regarding CD in Tab. 9, and our finetuned TensoSDF provides the best performance. Note that we do not present the results of optimizing full-resolution TensoSDF with Gaussians, since this variant always leads to out-of-memory on the Glossy dataset.
[…]
Angel Cat Horse
TSDF Fusion 0.63 2.22 0.61
Low-res. TensoSDF 0.61 1.35 0.44
finetuned TensoSDF 0.41 1.34 0.34
Table 10. Ablation study of three key components on the Glossy Blender dataset. “Def.” means the deferred Gaussian splatting pipeline, “Inc.” means the incorporation of SDF and Gaussian, and “Pru.” means the pruning operation.
@@ 6.4. Discussion and limitations
Figure 13. Due to the lack of indirect illumination modeling, our method fails to decompose reasonable material parameters (e.g., roughness) on the concave surface with complex inter-reflection.
In this part, we discuss several key differences between our method and other SDF-based methods. We mainly discuss two groups of closely related methods. The first group is the ones that optimize GS and SDF jointly (e.g., NeuSG, GSDF), and the second is the volumetric SDF methods. Then, we demonstrate some limitations of our methods and their potential solutions.
GS with SDF
Existing methods that optimize GS with SDF are designed for surface reconstruction and model the appearance by radiance, thus not supporting relighting. Due to the simplified appearance model, these methods fail to reconstruct reflective objects. NeuSG links SDF with GS only by aligning their normal. SDF and GS benefit marginally from each other, so it still takes 16 hours on an RTX 4090 for training. GSDF accelerates ray sampling of SDF by Gaussian depth and controls Gaussian opacity by SDF value. However, their method optimizes a full-resolution SDF with GS, which still takes 2 hours on an RTX A100 for training and encounters out-of-memory (OOM) issues. On the contrary, we utilize a low-resolution TensoSDF to regularize GS with mutual supervision, which is more efficient and avoids OOM. Additionally, we use the normal from fixed GS to enhance the geometry details of full-resolution TensoSDF efficiently. Note that though DeferredGS introduces SDF into GS, it only uses the pretrained SDF to regularize GS, while SDF cannot benefit from GS.
Volumetric SDF
Another group is the volumetric SDF-based methods that support relighting. As discussed in related works, these methods take a long time to converge, especially for the ones that handle reflective objects. We compare the open-source methods in the experiment section, while Neural-PBIR does not release the code. Its idea of introducing the differentiable path tracing to model complex inter-reflection could benefit the Gaussian-based methods.
Limitations
Our model improves the relighting and mesh quality compared to other Gaussian-based methods. However, some issues remain to be solved. We obverse the split-sum approximation in DiffRast (Laine et al., 2020) causes blurry renderings for low-roughness surfaces. We render a mirror-like plane using the Cycle path-tracing renderer in Blender and the split-sum approximation in the DiffRast. The comparison is in Fig. 12, and the split-sum result is blurry. We believe replacing it with more accurate rendering methods will improve the relighting quality. Last, we only consider direct lighting in our framework, which causes inconsistent material decomposition in the regions with complex inter-reflection, as shown in Fig. 13. Introducing indirect illumination will benefit the relighting quality. We leave it for future work.
Figure 14. Relighting results on the Glossy Blender Dataset. R3DG, GS-IR, TensoIR, and MII fail to model specular highlights. GShader provides specular relighting results with obvious artifacts. Our method, along with NeRO and TensoSDF, provides high-quality results, while our method only uses 25% training time of TensoSDF and enables real-time relighting. We present the metrics per scene (PSNR/SSIM).
Figure 15. Decomposed roughness visualization on the Glossy Blender Dataset. Gaussian-based methods and TensoIR reconstruct problematic roughness. MII provides over-smooth prediction and ignores the difference of material. Our method, NeRO, and TensoSDF reconstruct reasonable roughness for relighting, while our method takes 25% training time at most.
@@
Figure 18. Relighting results on real data from NeILF++. R3DG and GS-IR fail to reconstruct the reflective objects in real scenes. GShader cannot relight the specular highlights faithfully, while our method can provide reasonable relighting results.
@@
Figure 22. Mesh comparison on the Glossy Blender dataset. Our method provides high-quality meshes from the finetuned TensoSDF. All Gaussian-based methods fail to reconstruct smooth surfaces for reflective objects, while NeRF-based methods cannot preserve the geometry details. Our method ensures both global smoothness and local details. Note that the ground-truth meshes of this dataset are unavailable (only point cloud is available), so we choose to show the RGB reference under the nearest viewpoint on the dataset. Due to the limited space, we present the PGSR results in the supplementary.